# Constructions, bounds, and algorithms for peaceable queens

### v2

@@ Abstract.
Matthew Drescher was supported by the National Science Foundation under Grant #2127309 to the Computing Research Association for the CIFellows 2021 Project.
@@ 1. Introduction
The earliest known reference for the peaceable queens problem is actually Stephen Ainley’s 1977 book Mathematical Puzzles [1]. In his book, Ainley gives a construction which shows that $a(n)\geqslant\lfloor\frac{7}{48}n^{2}\rfloor$.
@@
$t(n)\geqslant 0.1339n^{2}$, for all sufficiently large even $n$.
@@
$t(n)\geqslant 0.0833n^{2}$, for all sufficiently large odd $n$.
@@
Our bound $a(n)\leqslant 0.1716n^{2}$ from Theorem 1.5 improves on the bound $a(n)\leqslant 0.25n^{2}$ proven by van Bommel and MacEachern [15]. The proof from [15] does not actually use the fact that queens can attack diagonally. In contrast, our non-linear program crucially exploits this fact. This is necessary since there is a lower bound of $0.25n^{2}$ for the ‘peaceable rooks’ problem. In principle, it may be possible to match the lower bound $\frac{7}{48}n^{2}\approx 0.1458n^{2}$ given by the Ainley construction by adding more valid constraints to our non-linear program.
@@
Paper Outline. We begin with some notation in Section 2. In Section 3, we describe our explicit constructions which prove Theorems 1.1, and 1.3. In Section 4, we define our non-linear programs and prove Theorem 1.2Theorem 1.4, and Theorem 1.5. We describe our algorithm in Section 5. In Appendix A we list the best odd torus solutions found by our algorithm.
@@ 3. Lower bounds
$B_{1}$ =$=\{(i,j)\subseteq[a]\times[b]:\text{$i+j$ is odd}\},$
$B_{2}$ =$=\{(i,j)\subseteq[a]\times[b+1,2n]:\text{$i$ is even and $j$ is odd}\},$
$B_{3}$ =$=\{(i,j)\subseteq[a+1,2n]\times[b]:\text{$i$ is odd and $j$ is even}\},\text{ and}$
$W$ =$=\{(i,j)\subseteq[a+1,2n]\times[b+1,2n]:\text{$i$ is even and $j$is even}\}.$
See Figure 2 for a drawing of the $(8,10)$-plaid on $\mathbb{T}_{32}$.
@@
$|B|\approx\frac{1}{2}ab+\frac{1}{4}a(n-b)+\frac{1}{4}b(n-a)\approx\frac{c^{2}n^{2}}{2}+\frac{c(1-c)n^{2}}{2}=\frac{c}{2}n^{2}.$
@@
$|W|\approx\frac{1}{4}(n-a)(n-b)\approx\frac{(1-c)^{2}}{4}n^{2}=\frac{c}{2}n^{2}.$
The above computations clearly hold up to an $O(n)$ error. Thus, $\min\{|B|,|W|\}\geqslant\frac{c}{2}n^{2}-O(n)$, as required. ∎
@@
Partition $[n]\times[n]$ into a set $\mathcal{T}$ of 36 congruent ‘triangles’ such that each of $B$ and $W$ intersects exactly 12 triangles. From Figure 3, it is clear what $\mathcal{T}$ should be in general. Let $T_{B}$ and $T_{W}$ be the union of the triangles which intersect $B$ and $W$ respectively. Note that $B$ covers roughly $\frac{1}{4}$ of the cells in $T_{B}$, and $W$ covers roughly $\frac{1}{4}$ of the cells in $T_{W}$. Thus,
$\min\{|B|,|W|\}\approx\frac{12}{36}\cdot\frac{1}{4}n^{2}=\frac{1}{12}n^{2}.$
Since the above argument clearly holds up to an $O(n)$ error, we are done. ∎
@@ 4.1. Odd torus
Let $n\in\mathbb{N}$ be odd and let $(B,W)$ be a largest peaceful battle on $\mathbb{T}_{n}$ with $|B|=|W|$. Suppose that $B$ intersects exactly $r$ rows, $c$ columns, $d$ $\mathbb{Z}_{n}$-diagonals and $a$ $\mathbb{Z}_{n}$-skew-diagonals of $\mathbb{T}_{n}$. Let $R,C,D,A\subseteq V(\mathbb{T}_{n})$ be the union of these rows, columns, $\mathbb{Z}_{n}$-diagonals, and $\mathbb{Z}_{n}$-skew-diagonals, respectively. Thus, $|R|=rn$, $|C|=cn$, $|D|=dn$ and $|A|=an$. Since $B$ is contained in $R\cap C\cap D\cap A$ and $W$ is disjoint from $R\cap C\cap D\cap A$, we have
$|R\cap C\cap D\cap A|\geqslant|B|\quad\text{ and }\quad|\overline{R\cup C\cup D\cup A}|\geqslant|W|.$ (1)
$y_{D}$$y_{R}$$y_{A}$$y_{C}$$z_{D\cap\overline{R}\cap\overline{A}\cap\overline{C}}$$z_{D\cap R\cap\overline{A}\cap\overline{C}}$$z_{D\cap C\cap\overline{R}\cap\overline{A}}$$z_{D\cap C\cap R\cap\overline{A}}$$z_{A\cap\overline{R}\cap\overline{D}\cap\overline{C}}$$z_{A\cap R\cap C\cap\overline{D}}$$z_{A\cap C\cap\overline{R}\cap\overline{D}}$$z_{A\cap R\cap\overline{D}\cap\overline{C}}$$z_{A\cap C\cap D\cap\overline{R}}$$z_{A\cap R\cap D\cap\overline{C}}$$z_{R\cap\overline{D}\cap\overline{A}\cap\overline{C}}$$z_{C\cap\overline{R}\cap\overline{A}\cap\overline{D}}$$z_{D\cap R\cap C\cap A}$
Figure 4. Each $y_{X}$ variable represents the fraction of vertices of $\mathbb{T}_{n}$ in $X$. Each $z_{S}$ variable represents the fraction of vertices of $\mathbb{T}_{n}$ contained in that region of the Venn Diagram.
Observe that each hyperedge of $\mathbb{T}_{n}$ has $n$ vertices and if $e$ and $e^{\prime}$ are distinct hyperedges of the same type, then $e\cap e^{\prime}=\emptyset$. On the other hand, if $e$ and $e^{\prime}$ are of different types, then $|e\cap e^{\prime}|=1$. Thus, the following equations hold.
$|X|$ $=xn\qquad\text{ for all }(x,X)\in\{(r,R),(c,C),(d,D),(a,A)\}$ (2)
$|X\cap X^{\prime}|$ $=xx^{\prime}\qquad\text{for all distinct }(x,X),(x^{\prime},X^{\prime})\in\{(r,R),(c,C),(d,D),(a,A)\}$ (3)
We now introduce variables to encode the above equations. Let $\mathcal{S}$ be the set of 16 regions of the Venn diagram of $R,C,D$, and $A$. For each $S\in\mathcal{S}$, we let $z_{S}:=\frac{|S|}{n^{2}}\in[0,1]$. See Figure 4 for a portion of the corresponding Venn diagram.
@@
$\sum_{S\in\mathcal{S}}z_{S}$ $=1$
$y_{X}$ $=\sum_{\begin{subarray}{c}S\in\mathcal{S}\\ S\subseteq X\end{subarray}}z_{S},\qquad$ for all $X\in\{R,C,D,A\}$
$y_{X\cap X^{\prime}}$ $=\sum_{\begin{subarray}{c}S\in\mathcal{S}\\ S\subseteq X\cap X^{\prime}\end{subarray}}z_{S},\qquad$ for all $X,X^{\prime}\in\{R,C,D,A\}$
$y_{X\cap X^{\prime}}$ $=y_{X}y_{X^{\prime}},\qquad$ for all distinct $X,X^{\prime}\in\{R,C,D,A\}$
$z_{S}$ $\in[0,1],\qquad$ for all $S\in\mathcal{S}$
@@ 4.2. Even torus
We now proceed as in the odd torus, except that there are more variables. To be precise, suppose that $B$ intersects exactly $r$ rows, $c$ columns, $d_{0}$ even $\mathbb{Z}_{n}$-diagonals, $d_{1}$ odd $\mathbb{Z}_{n}$-diagonals , $a_{0}$ even $\mathbb{Z}_{n}$-skew-diagonals, and $a_{1}$ odd $\mathbb{Z}_{n}$-skew-diagonals of $\mathbb{T}_{n}$. Let $R,C,D_{0},D_{1},A_{0},A_{1}\subseteq V(\mathbb{T}_{n})$ be the union of these rows, columns, even $\mathbb{Z}_{n}$-diagonals, odd $\mathbb{Z}_{n}$-diagonals, even $\mathbb{Z}_{n}$-skew-diagonals, and odd $\mathbb{Z}_{n}$-skew-diagonals, respectively. Thus, $|R|=rn$, $|C|=cn$, $|D_{0}|=d_{0}n$, $|D_{1}|=d_{1}n$, $|A_{0}|=a_{0}n$, and $|A_{1}|=a_{1}n$. Observe that
$|R\cap C\cap D_{0}\cap A_{0}|+|R\cap C\cap D_{1}\cap A_{1}|\geqslant|B|\quad\text{ and }\quad|\overline{R\cup C\cup D_{0}\cup D_{1}\cup A_{0}\cup A_{1}}|\geqslant|W|.$
Let $e\in E(\mathbb{T}_{n})$ and $e^{\prime}\in E(\mathbb{T}_{n})$. If $e$ is a row and $e^{\prime}$ is a column, then $|e\cap e^{\prime}|=1$. If $e$ is an even (respectively, odd) $\mathbb{Z}_{n}$-diagonal and $e^{\prime}$ is an even (respectively, odd) $\mathbb{Z}_{n}$-skew-diagonal, then $|e\cap e^{\prime}|=2$. If $e$ is a row or column and $e^{\prime}$ is a $\mathbb{Z}_{n}$-diagonal or a $\mathbb{Z}_{n}$-skew-diagonal, then $|e\cap e^{\prime}|=1$. Finally, if $e$ is an odd $\mathbb{Z}_{n}$-diagonal (respectively, odd $\mathbb{Z}_{n}$-skew-diagonal) and $e^{\prime}$ is an even $\mathbb{Z}_{n}$-diagonal (respectively, even $\mathbb{Z}_{n}$-skew-diagonal), then $|e\cap e^{\prime}|=0$. Thus, we have the following equations.
[…]
$|R\cap C|$ $=rc$
$|D_{0}\cap A_{0}|$ $=2d_{0}a_{0}$
$|D_{1}\cap A_{1}|$ $=2d_{1}a_{1}$
$|X\cap X^{\prime}|$ $=xx^{\prime}$ if $(x,X)\in\{(r,R),(c,C)\}$ and $(x^{\prime},X^{\prime})\in\{(d_{0},D_{0}),(d_{1},D_{1}),(a_{0},A_{0}),(a_{1},A_{1})\}$
$|X\cap X^{\prime}|$ $=0$ if $(x,X)\in\{(d_{0},D_{0}),(a_{0},A_{0})\}$ and $(x^{\prime},X^{\prime})\in\{(d_{1},D_{1}),(a_{1},A_{1})\}$
@@
$\sum_{S\in\mathcal{S}}z_{S}$ $=1$
$y_{X}$ $:=\sum_{\begin{subarray}{c}S\in\mathcal{S}\\ S\subseteq X\end{subarray}}z_{S},$ for all $X\in\boxtimes$
$y_{X\cap X^{\prime}}$ $:=\sum_{\begin{subarray}{c}S\in\mathcal{S}\\ S\subseteq X\cap X^{\prime}\end{subarray}}z_{S},$ for all $X,X^{\prime}\in\boxtimes$
@@
$y_{D_{0}\cap A_{0}}$ $=2y_{D_{0}}y_{A_{0}}$
$y_{D_{1}\cap A_{1}}$ $=2y_{D_{1}}y_{A_{1}}$
$y_{X\cap X^{\prime}}$ $=y_{X^{\prime}}y_{X},$ if $X\in\mathbin{\square}$ and $X^{\prime}\in\times$
$y_{X\cap X^{\prime}}$ $=0,$ if $X\in\{D_{0},A_{0}\}$ and $X^{\prime}\in\{D_{1},A_{1}\}$
$y_{X}$ $\in[0,\frac{1}{2}],$ for all $X\in\times$
$z_{S}$ $\in[0,1],$ for all $S\in\mathcal{S}$
@@ 4.3. Regular Board
Let $\mathcal{S}$ be the set of 64 regions of the Venn diagram of $R,C,D_{0},D_{1},A_{0}$, and $A_{1}$. For each $S\in\mathcal{S}$, we let $z_{S}:=\frac{|S|}{n^{2}}\in[0,1]$. Again we have
$|R\cap C\cap D_{0}\cap A_{0}|+|R\cap C\cap D_{1}\cap A_{1}|\geqslant|B|\quad\text{ and }\quad|\overline{R\cup C\cup D_{0}\cup D_{1}\cup A_{0}\cup A_{1}}|\geqslant|W|.$
@@
Suppose $\lfloor\alpha n\rfloor$ is odd (the case that $\lfloor\alpha n\rfloor$ is even is similar). Let $X\in\mathbin{\square}$, $X^{\prime}\in\times$ with $|X|\leqslant\alpha n^{2}$ and $|X^{\prime}|\leqslant\beta n^{2}$. Let $E_{X^{\prime}}$ be the hyperedges of type $X^{\prime}$ contained in $X^{\prime}$. For each $e\in E_{X^{\prime}}$ define $r(e):=\frac{|e\cap X|}{|e|}$. Sort $E_{X^{\prime}}$ in decreasing order according to $r(e)$. Let $\ell$ be the number of hyperedges of $E_{X^{\prime}}$ with $r(e)=1$. Observe that $\ell\leqslant\lceil\frac{\alpha n}{2}\rceil$. Moreover, for each $i\in\mathbb{N}$, the $(\ell+i)$th hyperedge in this ordering satisfies $r(e)\leqslant\frac{\lfloor\alpha n\rfloor}{\lfloor\alpha n\rfloor+2i}$. Let $Y$ consist of the first $\lfloor\alpha n\rfloor$ rows, and $Y^{\prime}$ consist of all cells $(i,j)$ such that $j-i\in\{0,2,4,\dots,2k\}$, where $k$ is the largest integer strictly less than $\lfloor\sqrt{\beta}n\rfloor$. Note that $|Y|=n\lfloor\alpha n\rfloor\leqslant\alpha n^{2}$ and $|Y^{\prime}|\leqslant\frac{(2k)^{2}}{4}\leqslant\frac{(2\sqrt{\beta}n)^{2}}{4}=\beta n^{2}$. Let $E_{Y^{\prime}}$ be the even anti-diagonals contained in $Y^{\prime}$. For each $e\in E_{Y^{\prime}}$ define $r(e):=\frac{|e\cap Y|}{|e|}$. Sort $E_{Y^{\prime}}$ in decreasing order according to $r(e)$. Let $\ell^{\prime}$ be the number of hyperedges of $E_{Y^{\prime}}$ with $r(e)=1$. Note that $\ell^{\prime}=\lceil\frac{\alpha n}{2}\rceil\geqslant\ell$. Moreover, for each $i\in\mathbb{N}$, the $(\ell+i)$th hyperedge in this ordering satisfies $r(e)=\frac{\lfloor\alpha n\rfloor}{\lfloor\alpha n\rfloor+2i}$. Therefore, $|Y\cap Y^{\prime}|\geqslant|X\cap X^{\prime}|$. We conclude the proof by observing that
$|Y\cap Y^{\prime}|\approx\frac{(\alpha n)(2\sqrt{\beta}n)}{2}-\frac{(\alpha n)^{2}}{4}=\alpha\sqrt{\beta}n^{2}-\frac{\alpha^{2}}{4}n^{2},$
where the above approximation clearly holds up to a $O(n)$ error. ∎
Let $S_{1}:=\overline{R\cup C\cup D\cup A}$, $S_{2}:=R\cap C\cap D_{0}\cap\overline{D_{1}}\cap A_{0}\cap\overline{A_{1}}$, and $S_{3}:=R\cap C\cap\overline{D_{0}}\cap D_{1}\cap\overline{A_{0}}\cap A_{1}$. Consider the following optimization problem.
@@
$\sum_{S\in\mathcal{S}}z_{S}$ $=1$
$y_{X}$ $=\sum_{\begin{subarray}{c}S\in\mathcal{S}\\ S\subseteq X\end{subarray}}z_{S},$ for all $X\in\boxtimes$
$y_{X\cap X^{\prime}}$ $=\sum_{\begin{subarray}{c}S\in\mathcal{S}\\ S\subseteq X\cap X^{\prime}\end{subarray}}z_{S},$ for all $X,X^{\prime}\in\boxtimes$
@@
$y_{X}^{2}+4y_{X\cap X^{\prime}}$ $\leqslant 4y_{X}\sqrt{y_{X^{\prime}}},$ if $X\in\mathbin{\square}$, $X^{\prime}\in\times$, $y_{X}\leqslant 2\sqrt{y_{X^{\prime}}}\leqslant 1$
$y_{X}$ $\in[0,\frac{1}{2}],$ for all $X\in\times$
$z_{S}$ $\in[0,1],$ for all $S\in\mathcal{S}$

### v3

@@ 1. Introduction
The earliest known reference for the peaceable queens problem is actually Stephen Ainley’s 1977 book Mathematical Puzzles [1]. In his book, Ainley gives a construction which shows that $a(n)\geqslant\lfloor\frac{7}{48}n^{2}\rfloor\approx 0.1458n^{2}$.
@@
$t(n)\geqslant 0.1339n^{2},$
for all sufficiently large even $n$.
@@
$t(n)\geqslant 0.0833n^{2},$
for all sufficiently large odd $n$.
@@
Our bound $a(n)\leqslant 0.1716n^{2}$ from Theorem 1.5 improves on the bound $a(n)\leqslant 0.25n^{2}$ proven by van Bommel and MacEachern [15]. The proof from [15] does not actually use the fact that queens can attack diagonally. In contrast, our non-linear program crucially exploits this fact. This is necessary since there is a lower bound of $0.25n^{2}$ for the ‘peaceable rooks’ problem. In principle, it may be possible to match the lower bound $\approx 0.1458n^{2}$ given by the Ainley construction by adding more valid constraints to our non-linear program.
@@
Paper Outline. We begin with some notation in Section 2. In Section 3, we describe our explicit constructions which prove Theorems 1.1, and 1.3. In Section 4, we define our non-linear programs and prove Theorem 1.2, Theorem 1.4, and Theorem 1.5. We describe our algorithm in Section 5. In Appendix A we list the best odd torus solutions found by our algorithm.
@@ 3. Lower bounds
$B_{1}$ =$=\{(i,j)\in[a]\times[b]:\text{$i+j$ is odd}\},$
$B_{2}$ =$=\{(i,j)\in[a]\times[b+1,2n]:\text{$i$ is even and $j$ is odd}\},$
$B_{3}$ =$=\{(i,j)\in[a+1,2n]\times[b]:\text{$i$ is odd and $j$ is even}\},$
and $W$ is the set of all pairs $(i,j)\in[a+1,2n]\times[b+1,2n]$ with $i$ and $j$ both even. See Figure 2 for a drawing of the $(8,10)$-plaid on $\mathbb{T}_{32}$.
@@
$|B|$ $\approx\frac{1}{2}ab+\frac{1}{4}a(n-b)+\frac{1}{4}b(n-a)$
$\approx\frac{c^{2}n^{2}}{2}+\frac{c(1-c)n^{2}}{2}$
$=\frac{c}{2}n^{2}.$
@@
$|W|\approx\frac{1}{4}(n-a)(n-b)\approx\frac{(1-c)^{2}}{4}n^{2}=\frac{c}{2}n^{2},$
where the last equality is the only place in the proof where we use the explicit value of $c$. Moreover, the above computations clearly hold up to an $O(n)$ error. Thus, $\min\{|B|,|W|\}\geqslant\frac{c}{2}n^{2}-O(n)$, as required. ∎
@@
We cut a square $S$ of side length $n$ into a set of 36 congruent triangles $T_{1},\dots,T_{36}$ as follows. First cut $S$ into 9 congruent squares, and then cut each of the smaller 9 squares into four congruent triangles by cutting along their two diagonals. Next, we regard each cell $(i,j)\in[n]\times[n]$ as a $1\times 1$ square contained in $S$.
Figure 4. Partition of $[n]\times[n]$ into $36$ sets.
See Figure 4 for a picture of the $33$-argyle embedded inside a square of side length $33$. For each cell $x$, we choose $f(x)\in[36]$ such that $x\cap T_{f(x)}\neq\emptyset$. Note that the choice of $f(x)$ is unique for all but at most $10n$ cells of $[n]\times[n]$. Let $T_{1}^{\prime},\dots,T_{36}^{\prime}$ be the partition of $[n]\times[n]$, where $x\in T_{f(x)}^{\prime}$ for all $x\in[n]\times[n]$. From Figure 4, it is clear that up to an $O(n)$ error, $B$ intersects 12 of $T_{1}^{\prime},\dots,T_{36}^{\prime}$ with density $\frac{1}{4}$ and 24 of $T_{1}^{\prime},\dots,T_{36}^{\prime}$ with density 0. The same holds for $W$. Thus, up to an $O(n)$ error we have $\min\{|B|,|W|\}\approx\frac{12}{36}\cdot\frac{1}{4}n^{2}=\frac{1}{12}n^{2}.$
@@ 4.1. Odd torus
Let $n\in\mathbb{N}$ be odd and let $(B,W)$ be a largest peaceful battle on $\mathbb{T}_{n}$ with $|B|=|W|$. Suppose that $B$ intersects exactly $r$ rows, $c$ columns, $d$ $\mathbb{Z}_{n}$-diagonals and $s$ $\mathbb{Z}_{n}$-skew-diagonals of $\mathbb{T}_{n}$. Let $R,C,D,S\subseteq V(\mathbb{T}_{n})$ be the union of these rows, columns, $\mathbb{Z}_{n}$-diagonals, and $\mathbb{Z}_{n}$-skew-diagonals, respectively. Thus, $|R|=rn$, $|C|=cn$, $|D|=dn$ and $|S|=sn$. Since $B$ is contained in $R\cap C\cap D\cap S$ and $W$ is disjoint from $R\cap C\cap D\cap S$, we have $|R\cap C\cap D\cap S|\geqslant|B|$ and $|\overline{R\cup C\cup D\cup S}|\geqslant|W|$.
Observe that each hyperedge of $\mathbb{T}_{n}$ has $n$ vertices and if $e$ and $e^{\prime}$ are distinct hyperedges of the same type, then $e\cap e^{\prime}=\emptyset$. On the other hand, if $e$ and $e^{\prime}$ are of different types, then $|e\cap e^{\prime}|=1$. Thus, for all distinct $(x,X),(x^{\prime},X^{\prime})\in\{(r,R),(c,C),(d,D),(s,S)\}$ we have $|X\cap X^{\prime}|=xx^{\prime}$.
We now introduce variables to encode the above equations. Let $\mathcal{F}$ be the set of 16 regions of the Venn diagram of $R,C,D$, and $S$. For each $F\in\mathcal{F}$, we let $z_{F}:=\frac{|F|}{n^{2}}\in[0,1]$. See Figure 5 for a depiction [^2] Note that some $z_{F}$ variables are missing from Figure 5 since it is impossible to draw 16 non-empty regions if we use disks to represent the four sets. of some of the variables $z_{F}$.
$y_{D}$$y_{R}$$y_{S}$$y_{C}$$z_{D\cap\overline{R}\cap\overline{S}\cap\overline{C}}$$z_{D\cap R\cap\overline{S}\cap\overline{C}}$$z_{D\cap C\cap\overline{R}\cap\overline{S}}$$z_{D\cap C\cap R\cap\overline{S}}$$z_{S\cap\overline{R}\cap\overline{D}\cap\overline{C}}$$z_{S\cap R\cap C\cap\overline{D}}$$z_{S\cap C\cap\overline{R}\cap\overline{D}}$$z_{S\cap R\cap\overline{D}\cap\overline{C}}$$z_{S\cap C\cap D\cap\overline{R}}$$z_{S\cap R\cap D\cap\overline{C}}$$z_{R\cap\overline{D}\cap\overline{S}\cap\overline{C}}$$z_{C\cap\overline{R}\cap\overline{S}\cap\overline{D}}$$z_{D\cap R\cap C\cap S}$
Figure 5. Each $y_{X}$ variable represents the fraction of vertices of $\mathbb{T}_{n}$ in $X$. Each $z_{F}$ variable represents the fraction of vertices of $\mathbb{T}_{n}$ contained in that region of the Venn Diagram.
@@
$\sum_{F\in\mathcal{F}}z_{F}=1$
$y_{X}=\sum_{\begin{subarray}{c}F\in\mathcal{F}\\ F\subseteq X\end{subarray}}z_{F},\,\text{for all $X\in\{R,C,D,S\}$}$
$y_{X\cap X^{\prime}}=\sum_{\begin{subarray}{c}F\in\mathcal{F}\\ F\subseteq X\cap X^{\prime}\end{subarray}}z_{F},\,\text{for all $X,X^{\prime}\in\{R,C,D,S\}$}$
$y_{X\cap X^{\prime}}=y_{X}y_{X^{\prime}},\,\text{for all distinct $X,X^{\prime}\in\{R,C,D,S\}$}$
$z_{F}\in[0,1],\,\text{for all $F\in\mathcal{F}$}$
@@ 4.2. Even torus
We now proceed as in the odd torus, except that there are more variables. To be precise, suppose that $B$ intersects exactly $r$ rows, $c$ columns, $d_{0}$ even $\mathbb{Z}_{n}$-diagonals, $d_{1}$ odd $\mathbb{Z}_{n}$-diagonals, $s_{0}$ even $\mathbb{Z}_{n}$-skew-diagonals, and $s_{1}$ odd $\mathbb{Z}_{n}$-skew-diagonals of $\mathbb{T}_{n}$. Let $R,C,D_{0},D_{1},S_{0},S_{1}\subseteq V(\mathbb{T}_{n})$ be the union of these rows, columns, even $\mathbb{Z}_{n}$-diagonals, odd $\mathbb{Z}_{n}$-diagonals, even $\mathbb{Z}_{n}$-skew-diagonals, and odd $\mathbb{Z}_{n}$-skew-diagonals, respectively. Thus, $|R|=rn$, $|C|=cn$, $|D_{0}|=d_{0}n$, $|D_{1}|=d_{1}n$, $|S_{0}|=s_{0}n$, and $|S_{1}|=s_{1}n$. Observe that
$|R\cap C\cap D_{0}\cap S_{0}|+|R\cap C\cap D_{1}\cap S_{1}|\geqslant|B|,$
and
$|\overline{R\cup C\cup D_{0}\cup D_{1}\cup S_{0}\cup S_{1}}|\geqslant|W|.$
Let $e\in E(\mathbb{T}_{n})$ and $e^{\prime}\in E(\mathbb{T}_{n})$. If $e$ is a row and $e^{\prime}$ is a column, then $|e\cap e^{\prime}|=1$. If $e$ is an even (respectively, odd) $\mathbb{Z}_{n}$-diagonal and $e^{\prime}$ is an even (respectively, odd) $\mathbb{Z}_{n}$-skew-diagonal, then $|e\cap e^{\prime}|=2$. If $e$ is a row or column and $e^{\prime}$ is a $\mathbb{Z}_{n}$-diagonal or a $\mathbb{Z}_{n}$-skew-diagonal, then $|e\cap e^{\prime}|=1$. Finally, if $e$ is an odd $\mathbb{Z}_{n}$-diagonal (respectively, odd $\mathbb{Z}_{n}$-skew-diagonal) and $e^{\prime}$ is an even $\mathbb{Z}_{n}$-diagonal (respectively, even $\mathbb{Z}_{n}$-skew-diagonal), then $|e\cap e^{\prime}|=0$. Thus, $|R\cap C|=rc$, $|D_{0}\cap S_{0}|=2d_{0}s_{0}$, $|D_{1}\cap S_{1}|=2d_{1}s_{1}$; and $|X\cap X^{\prime}|=0$ if $(x,X)\in\{D_{0},S_{0}\}$ and $X^{\prime}\in\{D_{1},S_{1}\}$; and $|X\cap X^{\prime}|=xx^{\prime}$ if $(x,X)\in\{(r,R),(c,C)\}$ and $(x^{\prime},X^{\prime})\in\{(d_{0},D_{0}),(d_{1},D_{1}),(s_{0},S_{0}),(s_{1},S_{1})\}$.
@@
$\sum_{F\in\mathcal{F}}z_{F}=1$
$y_{X}:=\sum_{\begin{subarray}{c}F\in\mathcal{F}\\ F\subseteq X\end{subarray}}z_{F},\text{ for all $X\in\boxtimes$}$
$y_{X\cap X^{\prime}}:=\sum_{\begin{subarray}{c}F\in\mathcal{F}\\ F\subseteq X\cap X^{\prime}\end{subarray}}z_{F},\text{ for all $X,X^{\prime}\in\boxtimes$}$
@@
$y_{D_{0}\cap S_{0}}=2y_{D_{0}}y_{S_{0}}$
$y_{D_{1}\cap S_{1}}=2y_{D_{1}}y_{S_{1}}$
$y_{X\cap X^{\prime}}=y_{X^{\prime}}y_{X},\text{ if $X\in\mathbin{\square}$ and $X^{\prime}\in\times$}$
$y_{X\cap X^{\prime}}=0,\text{ if $X\in\{D_{0},S_{0}\}$ and $X^{\prime}\in\{D_{1},S_{1}\}$}$
$y_{X}\in[0,\frac{1}{2}],\text{ for all $X\in\times$}$
$z_{F}\in[0,1],\text{ for all $F\in\mathcal{F}$}$
@@ 4.3. Regular Board
Let $\mathcal{F}$ be the set of 64 regions of the Venn diagram of $R,C,D_{0},D_{1},S_{0}$, and $S_{1}$. For each $F\in\mathcal{F}$, we let $z_{F}:=\frac{|F|}{n^{2}}\in[0,1]$. Again we have
$|R\cap C\cap D_{0}\cap S_{0}|+|R\cap C\cap D_{1}\cap S_{1}|\geqslant|B|,$
and
$|\overline{R\cup C\cup D_{0}\cup D_{1}\cup S_{0}\cup S_{1}}|\geqslant|W|.$
@@
Suppose $\lfloor\alpha n\rfloor$ is odd (the case that $\lfloor\alpha n\rfloor$ is even is similar). Let $X\in\mathbin{\square}$, $X^{\prime}\in\times$ with $|X|\leqslant\alpha n^{2}$ and $|X^{\prime}|\leqslant\beta n^{2}$. Let $E_{X^{\prime}}$ be the hyperedges of type $X^{\prime}$ contained in $X^{\prime}$. For each $e\in E_{X^{\prime}}$ define $r(e):=\frac{|e\cap X|}{|e|}$. Sort $E_{X^{\prime}}$ in decreasing order according to $r(e)$. Let $\ell$ be the number of hyperedges of $E_{X^{\prime}}$ with $r(e)=1$. Observe that $\ell\leqslant\lceil\frac{\alpha n}{2}\rceil$. Moreover, for each $i\in\mathbb{N}$, the $(\ell+i)$th hyperedge in this ordering satisfies $r(e)\leqslant\frac{\lfloor\alpha n\rfloor}{\lfloor\alpha n\rfloor+2i}$. Let $Y$ consist of the first $\lfloor\alpha n\rfloor$ rows, and $Y^{\prime}$ consist of all cells $(i,j)$ such that $j-i\in\{0,2,4,\dots,2k\}$, where $k$ is the largest integer strictly less than $\lfloor\sqrt{\beta}n\rfloor$. Note that $|Y|=n\lfloor\alpha n\rfloor\leqslant\alpha n^{2}$ and $|Y^{\prime}|\leqslant\frac{(2k)^{2}}{4}\leqslant\frac{(2\sqrt{\beta}n)^{2}}{4}=\beta n^{2}$. Let $E_{Y^{\prime}}$ be the even anti-diagonals contained in $Y^{\prime}$. For each $e\in E_{Y^{\prime}}$ define $r(e):=\frac{|e\cap Y|}{|e|}$. Sort $E_{Y^{\prime}}$ in decreasing order according to $r(e)$. Let $\ell^{\prime}$ be the number of hyperedges of $E_{Y^{\prime}}$ with $r(e)=1$. Note that $\ell^{\prime}=\lceil\frac{\alpha n}{2}\rceil\geqslant\ell$. Moreover, for each $i\in\mathbb{N}$, the $(\ell+i)$th hyperedge in this ordering satisfies $r(e)=\frac{\lfloor\alpha n\rfloor}{\lfloor\alpha n\rfloor+2i}$. Therefore, $|Y\cap Y^{\prime}|\geqslant|X\cap X^{\prime}|$. We conclude the proof by observing that $|Y\cap Y^{\prime}|\approx\frac{(\alpha n)(2\sqrt{\beta}n)}{2}-\frac{(\alpha n)^{2}}{4}=\alpha\sqrt{\beta}n^{2}-\frac{\alpha^{2}}{4}n^{2},$ where the above approximation clearly holds up to a $O(n)$ error. ∎
Let $F_{1}:=\overline{R\cup C\cup D\cup A}$, $F_{2}:=R\cap C\cap D_{0}\cap\overline{D_{1}}\cap S_{0}\cap\overline{S_{1}}$, and $F_{3}:=R\cap C\cap\overline{D_{0}}\cap D_{1}\cap\overline{S_{0}}\cap S_{1}$. Consider the following optimization problem.
@@
$\sum_{F\in\mathcal{F}}z_{F}=1$
$y_{X}=\sum_{\begin{subarray}{c}F\in\mathcal{F}\\ F\subseteq X\end{subarray}}z_{F},\text{ for all $X\in\boxtimes$}$
$y_{X\cap X^{\prime}}=\sum_{\begin{subarray}{c}F\in\mathcal{F}\\ F\subseteq X\cap X^{\prime}\end{subarray}}z_{F},\text{ for all $X,X^{\prime}\in\boxtimes$}$
@@
$y_{X}^{2}+4y_{X\cap X^{\prime}}\leqslant 4y_{X}\sqrt{y_{X^{\prime}}},$
if $X\in\mathbin{\square}$, $X^{\prime}\in\times$, $y_{X}\leqslant 2\sqrt{y_{X^{\prime}}}\leqslant 1$
$y_{X}\in[0,\frac{1}{2}],\text{ for all $X\in\times$}$
$z_{F}\in[0,1],\text{ for all $F\in\mathcal{F}$}$